A set of normally distributed data has mean 47 s and standard deviation 0.6 s. Find the percentage of the data within these intervals:
a. Between 46 and 47 s
b. Greater than 1.5s above or below the mean

Answer by stanbon(75887)

A set of normally distributed data has mean 47 and standard deviation 0.6 s. Find the percentage of the data within these intervals:
a. Between 46 and 47
Draw a normal curve with mean = x = 47
Plot 46 on the x-axis
Find the z-value of 47
z(46) = (46-47)/0.6 = -1/0.6 = -5/3
Note: The z-value of the mean is 0
-----
Percentage between 46 and 47 = P(-5/3 < z < 0) = 0.4522
-----------------------------------------------------
b. Greater than 1.5s above or below the mean
If that "s" means standard deviation the answer is:
normalcdf(-100,-1.5) = 0.0668
================================
Cheers,
Stan H.
==================